For n∈Z, prove n^2 is odd if and only if n is odd.


1
Expert's answer
2021-09-26T18:03:04-0400

For "n\in\Z", let us prove that "n^2" is odd if and only if "n" is odd.

Let "n" be an odd number. Then "n=2k+1" for some "k\in\Z." Then "n^2=(2k+1)^2=4k^2+4k+1=2(2k^2+2k)+1=2s+1," where "s=2k^2+2k\in\Z." Therefore, "n^2" is odd.

Let "n^2" be an odd number. Let us prove that "n" is odd using the method by contradiction. Suppose that "n" is not odd, and hence "n" is even, that is "n=2k" for some "k\in\Z." Then "n^2=(2k)^2=4k^2=2(2k^2)=2m," where "m=2k^2\in\Z." It follows that "n^2" is even. This contradiction proves that "n" is odd.